### v1

# Breakdown of sequential tunnel ionization in ultrashort electromagnetic pulses

Abstract
We consider double ionization of negative bromine ion ${\rm Br}^{-}$ in intense linearly polarized low-frequency laser fields. By solving numerically the two-electron time-dependent Schrödinger equation we demonstrate that while for electromagnetic pulses of a few tens of femtoseconds duration and longer the sequential single-electron approximation perfectly describes the ionization dynamics, for pulses as short as a few femtoseconds this picture breaks down entirely, and the electron-electron interaction suppresses the rate of ionization by approximately one order of magnitude. We also show weak but clear signatures of the collective tunneling effect in the photoelectron density distribution.
Tunnel ionization, electron-electron correlations, intense laser fields, collective tunneling
The fundamental effect of quantum mechanical tunneling discussed for the first time in the context of $\alpha$-decays [1, 2] is pivotal in the physics of nuclei [3], the physics of atoms in external fields [4, 5], and the physics of condensed matter [6]. Optical tunneling in atoms, molecules, and solids subject to intense infrared laser fields is the key phenomenon of contemporary nonlinear optics. It underlies attosecond science [7, 8], the generation of high-order harmonics (HHG) [9, 10], and attosecond spectroscopy [11, 12, 13]. Tunnel and multiphoton ionization of atoms in the field of intense laser radiation first theoretically addressed by Keldysh [14], was observed initially as multiquantum absorption [15] in the mid-60es, and later – in the form of above-threshold ionization (ATI) [16], a process where the electron absorbs more photons when necessary to escape from the atom. Both ATI and HHG can be efficiently explained within the semi-classical three-step model [17, 18], where optical tunneling serves as the first step.
A wide range of theoretical approaches describing the phenomenon of strong-field ionization and related processes are based on the single-active electron (SAE) approximation. This is commonly substantiated by the fact that, in the regime of optical tunneling, the electron must travel a sufficient distance from the nucleus to appear in that part of space where the laser force dominates over the Coulomb attraction. The higher the ionization potential $I_{p}$ of the level, the smaller is the probability that the electron appears at large distances from the nucleus, so that the model in which the currently outermost electron solely interacts with the laser field sounds adequate. We will refer to the regime of ionization dominated by the single-electron dynamics as sequential ionization. The celebrated exclusion from this rule takes place in linearly polarized fields of non-relativistic intensity, where the outermost electron may undergo a recollision with the atomic residual along the same scenario, as underlies the effects of HHG [18] and high-energy ATI plateau [19]. Recollision leads to an extreme enhancement in double ionization that is up to several orders of magnitude, compared to that expected on the basis of the sequential mechanism. This effect of nonsequential double ionization (NSDI) is the eminent example of an ionization channel strongly mediated by electron-electron interaction in the regime of tunneling. By now, NSDI in atoms has been thoroughly studied both experimentally and in theory so that its recollision genesis is beyond doubt; see Ref.[20] for a review. Other nosequential ionization mechanisms including the shake-off model [21] and the collective tunneling model [22, 23, 24] have been theoretically discussed and numerically examined [25, 23] but not yet observed in experiments.
In linearly polarized fields of nonrelativistic intensity, the NSDI through recollision dominates other possible ionization channels by orders of magnitude, which makes a search of alternative nonsequential mechanisms extremely difficult both in numerical calculations and in experiment. However, there are three ionization regimes where such alternatives can be revealed: (i) relativistic ionization, (ii) ionization in fields with circular or elliptical polarization where rescattering is damped out be the two-dimensional motion of the laser-driven electron (see e.g. [26, 27] for details and examples), and (iii) ionization by extremely short laser pulses with duration insufficient to bring the electron back to its parent ion. The latter is hardly achievable in experiment, but easy to reach in theoretical consideration, where modeling of strong-field ionization by short unipolar pulses allows for addressing tiny effects otherwise shielded by the oscillating electron’s dynamics [28].
Here, some elaboration of the terminology is required to avoid ambiguities and misinterpretations in the what follows. We characterize ionization as sequential if the electrons escape from the atom one after another with some significant time interval. The electron-electron interaction may or may not play an important role in this process, so that the sequential ionization is not necessarily equivalent to ionization of model non-interacting electrons. The nonsequential ionization is apparently something opposite, so it is a process in which both electrons depart from the ion simultaneously or within a small time window. Note that in the recollision induced NSDI the first electron leaves the atom approximately one laser period earlier than the second one. However, this fact does not withdraw this channel from the present definition, because eventually the electrons depart simultaneously, after the recollision, while the preceding evolution of the first electron is a virtual process. Note that the term NSDI is equally applicable to different mechanisms of double ionization [29, 30, 23, 25]. Simultaneous emission can take place even in the case of non-interacting electrons or for interacting electrons, but without recollision and excitation of intermediate bound states. We will refer to such simultaneous emission as collective tunneling [22, 23].
In this paper, we examine the ionization dynamics of a two-electron atomic system subject to a strong time-dependent electric field under conditions that eliminate recollision. We choose the bromine negative ion as a target with two very different ionization potentials to provide conditions, where the collective channel of ionization [22, 23] can also potentially make a sizable contribution [24]. The Keldysh parameter $\gamma=\sqrt{2mI_{p}}\omega/eE_{0}$ [14] is kept about unity or smaller, so that the ionization proceeds through the tunneling mechanism. Here $\omega$ and $E_{0}$ are the characteristic frequency and strength of the electric field, $I_{p}$ is the ionization potential, $e$ and $m$ are the elementary charge and mass correspondingly. Solving numerically the time-dependent Schrödinger equation (TDSE), we calculate the probability of double ionization for the 2-dimensional (2D) ${\rm Br}^{-}$ ion and show that, very surprisingly, in short pulses with duration $\approx 2$ fs this probability deviates dramatically from that expected in the case of a sequential process, when the electrons are being removed one after another. Not less surprisingly, in longer pulses of several dozens fs duration, a rudimentary sequential tunneling model reproduces the exact numerical result with accuracy close to one per cent. Thus in a single calculation we prove the validity of the sequential tunnel ionization picture for one set of parameters and demonstrate its complete failure for another. Not less importantly, we also show that the two-particle density distribution demonstrates a signature of the collective tunneling effect.
Our calculations made for a 2D two-electron atom, thus effectively for a 4D system, appear demanding as far as the required computer time and memory are concerned. For better justification of our findings we also present some results obtained for a 1D two-electron atom, which allow for tracing the wave function evolution at a much longer time scale.
We compare the ionization dynamics of two model systems: (a) a 2D two-electron negative ion with interacting electrons and (b) a 2D ion with two non-interacting electrons having the same ionization potentials as (a). In case (a), interaction with the parent ion $V_{\text{ion}}$ and electron-electron interaction $V_{\text{int}}$ are described by the smoothed Coulomb potentials (here and below we use atomic units $m=e=\hbar=1$)
$V_{\text{ion}}(\bm{r})=\frac{-1}{\sqrt{\bm{r}^{2}+a^{2}}}~{},$ (1)
$V_{\text{int}}(\bm{r}_{1},\bm{r}_{2})=\frac{1}{\sqrt{(\bm{r}_{1}-\bm{r}_{2})^{2}+b^{2}}}~{}.$ (2)
The smoothing parameter $a$ is chosen to reproduce the second ionization potential $I_{p2}$ when the outermost electron is removed. After this, the parameter $b$ is adjusted to reproduce the $I_{p1}$ of the outer electron. Parameters $a$ and $b$ and the ground-state wavefunction are found by numerically solving the TDSE (see details below).
The two-electron TDSE for the case (a) of the interacting electrons has the form
$\begin{gathered}i\frac{\partial}{\partial t}\Psi(\bm{r}_{1},\bm{r}_{2},t)=\\ =\big{[}H_{0}(\bm{r}_{1},t)+H_{0}(\bm{r}_{2},t)+V_{\text{int}}(\bm{r}_{1},\bm{r}_{2})\big{]}\Psi(\bm{r}_{1},\bm{r}_{2},t)~{},\end{gathered}$ (3)
where $H_{0}$ is the single particle Hamiltonian including interaction with the electric field of the laser pulse
$H_{0}(\bm{r},t)=-\frac{1}{2}\Delta\bm{r}+V_{\text{ion}}(\bm{r})-\varphi(\bm{r},t)~{}$ (4)
with $\varphi$ being its scalar potential.
To suppress recollisions, the electric field is set in the form of an unipolar pulse polarized along the $x$-axis with the amplitude value $E_{0}$ and duration $T$
$E(t)=-E_{0}\sin^{2}\left(\dfrac{\pi t}{T}\right)~{}.$ (5)
In this spatially homogeneous field the electrons are accelerated after ionization, rapidly reaching the boundaries of the computation box. This hinders analysis of the ionization dynamics. To make such analysis more affordable, in the numerical calculation we use the field which is homogeneous in the vicinity of the ion and vanishes far from it:
$\tilde{E}(x,t)=\begin{cases}E(t)\cos^{2}\left(\dfrac{\pi x}{2x_{0}}\right),~{}|x|\leq x_{0}\\ 0,~{}\text{elsewhere}~{},\end{cases}$ (6)
[…]
The TDSE (3) is solved for the 1D and 2D systems. For the 1D system coordinates $x_{1}$ and $x_{2}$ of the electrons were discretized on a spatial grid of $2048\times 2048$ nodes with a step $\Delta x=0.5$ a.u. The absorbing layer [31] occupies 15 nodes at the boundary of the computational box. The space envelope parameter in (6) was set to $x_{0}=50$ a.u. For the 2D problem, the coordinates were discretized with a step $\Delta x=\Delta y=0.5$ a.u. and $N_{x}=512$, $N_{y}=256$ nodes for each electron. The total size of the computational box is $512\times 256\times 512\times 256$ nodes. In this case, we took $x_{0}=30$ a.u. To reproduce the first $I_{p1}=3.37$ eV and second $I_{p2}=11.81$ eV ionization potentials of $\rm{Br}^{-}$ we used $a=1.66$, $b=2.6$ for 1D electrons and $a=1$, $b=2.2$ for 2D electrons in (1) and (2).
In the case (b), the TDSE with the single-particle Hamiltonian was solved for the first and second electrons separately. For the outer electron, the potential was approximated by a short-range well of the form
$V_{\text{short}}(\bm{r})=-c_{1}\exp\left(-\dfrac{\bm{r}^{2}}{c_{2}^{2}}\right)$ (7)
with two fitting parameters $c_{1}$ and $c_{2}$ chosen to reproduce the first ionization potential $I_{p1}$. For the second electron, we used the smoothed Coulomb potential (1) with a fitting parameter $a$ chosen to reproduce $I_{p2}$. These parameters were taken $c_{1}=0.48$, $c_{2}=0.76$ and $a=1.66$ for the 1D problem and $c_{1}=0.48$, $c_{2}=2$ and $a=1$ for the 2D problem.
In both cases (a) and (b) the initial state wave function $\Psi_{0}(\bm{r}_{1},\bm{r}_{2},t)$ was symmetrized with respect to coordinates assuming the full spin 0. The two-electron ground state wave function was obtained by using the imaginary time method [32]. The double ionization probability $W(t)$ was found by integrating the modulus squared wave function over space with the ion vicinity excluded
[…]
Here, $\varkappa=4..14$ is the size of the area around the ion, which was selected such that the probability $W$ does not depend on $t$ at the end of the calculation.
To represent the space probability distribution for two 2D electrons on the plane $(x_{1},x_{2})$ of the coordinates along the field direction, we integrate $|\Psi({\bf r}_{1},{\bf r}_{2},t)|^{2}$ over the transverse coordinates $y_{1}$ and $y_{2}$, excluding an area near the parent ion
[…]
Cutting out small values of $y_{1}$ and $y_{2}$ eliminates some part of the contribution from single ionization (when the second electron is located near $y=0$) and makes the double ionization flux better visible on the background generated by the outer electron.
We consider two ionization regimes. In the first case, we use a long electromagnetic pulse of amplitude $E_{0}=0.045$ (corresponding to the peak intensity $\approx 10^{14}{\rm W/cm}^{2}$), and a duration of $T=40$ fs. Below we refer this case to as the “long pulse regime” (LPR). In the LPR the ionization probability for the inner electron becomes sizable, when the outer orbital is already depleted. In this case, the electron-electron interaction is significant in the bound state only, and we expect the SAE approximation describing the ionization with reasonable accuracy. FIG. 1 shows the two-electron probability distribution for the interacting (a) and non-interacting (b) electrons in 1D. These distributions look rather similar, except for a small population near the diagonal $x_{1}=x_{2}$ present for the non-interacting electrons (b). The general similarity of this distributions indicates that, in the case of sufficiently long pulses, e-e interaction does not significantly affect the ionization process, because the most probable instants of detachment of the outer and the inner electrons are well separated in time. To provide a quantitative measure of this similarity, we integrate the populations over space to find the total probability of double ionization (8). It appears surprisingly insensitive to the presence of e-e interaction: the SAE approximation and the TDSE result (3) differ by less than $1\%$ (FIG. 2(a)). A similar calculation with the equally long pulse for 2D electrons is unaffordable in view of the requires computer resources, so we reduced the pulse duration to $T=13$ fs. The distribution function calculated for the 2D systems (9) is shown in panels (c) and (d). Qualitatively, they resemble the distributions of (a) and (b) with the only significant difference, which is the presence of a small diagonal population for the interacting electrons (c). The total probability of double ionization (8) for the interacting and non-interacting 2D electrons (FIG. 2(b)) differ more than in the 1D case (FIG. 2(a)), but we expect that the difference will decrease with the pulse duration increasing.
Figure 1: Probability density $|\Psi(x_{1},x_{2},t)|^{2}$ for the 1D ion (a,b) and distribution $F(x_{1},x_{2},t)$ calculated along (9) for the 2D ion (c,d) with the interacting (a,c) and non-interacting electrons (b,d). The time instants and the field parameters (different for the 1D and 2D cases) are: $t=17.3$ fs, $E_{0}=0.045$, $T=40$ fs (a,b) and $t=7.2$ fs, $E_{0}=0.045$, $T=13$ fs (c,d). The distributions are shown in relative units.
Figure 2: Total probability of double ionization (8) for the interacting (red lines) and non-interacting (blue lines) electrons in the 1D (a) and 2D (b) ion. The field parameters are: $E_{0}=0.045$, $T=40$ fs (a) and $T=13$ fs (b).
Now we turn to the criterion of separation between LPR and SPR. Fig.3 shows the populations $n_{1}$ of the single ionized species (neutral bromine) and $n_{2}$ of the doubly ionized species ${\rm Br}^{+}$ as functions of time. These results were obtained by numerically solving the system of rate equations
$\begin{cases}\dot{n_{0}}=-w_{1}n_{0}\\ \dot{n_{1}}=w_{1}n_{0}-w_{2}n_{1}\\ \dot{n_{2}}=w_{2}n_{1}~{},~{}~{}~{}n_{0}+n_{1}+n_{2}=1~{},~{}~{}~{}n_{0}(0)=1,\\ \end{cases}$ (10)
where $n_{0}$ is the population of the seeding specie (negative ion $\text{Br}^{-}$ in our case). For the time-dependent ionization rates $w_{1,2}[E(t)]$ in (10), we use the celebrated sequential tunnel Perelomov-Popov-Terent’ev (PPT) formulas [33, 34] (see also reviews [5, 35] for detailed descriptions). For the present case, we are not addressing effects of the bound state angular momentum on the rate of ionization, so $l=m=0$ was set in the PPT rates. Fig. 3 illustrates the qualitative difference between the LPR (a) and SPR (b): in the former case the outer orbital is entirely stripped out before the inner electron can be considerably affected by the laser field; in the SPR the inner electron begins to feel the field before the outer one is removed. The border between the two regimes is approximately determined by the ionization saturation time for the outer electron. Here we determine the saturation time as such pulse duration that the population $n_{1}$ of singly ionized atoms reaches 0.9 at the field maximum. Numerical analysis of (10) gives the saturation time for the outer electron $\approx 11$ fs for $E_{0}=0.045$ and $\approx 17$ fs for $E_{0}=0.035$. These numbers suggest that for the pulse duration $T\approx 2$ fs the SPR is amply achieved (Fig. 3(b)).
Figures 4 and 5 present the SPR analogues of Figs. 1 and 2, respectively. Their pairing comparison reveals both similarities and striking differences in the ionization dynamics. In Fig.4(b), a significant part of the electron density is concentrated near the diagonal $x_{1}=x_{2}$. This indicates that the non-interacting electrons tend to escape side-by-side much more than they do in the LPR for the reasons we explained above. Analysis of the distributions reveals two more effects present for the chosen SPR parameters: (i) considerable suppression of the two-electron ionization probability, when the e-e repulsion blocks the diagonal flux and (ii) simultaneous tunneling of two electrons. Firstly, comparison of the distributions shown in Fig.4(a) and (b) demonstrates that for the 1D ion the e-e interaction blocks the diagonal flux entirely, which must lead to a great decrease in double ionization, compared to the prediction of the SAE sequential model. The plots of Fig.5(a) show that this is indeed the case, and the total probability of double ionization drops by more than one order of magnitude. Secondly, in the more physically meaningful case of the 2D ion the effect of suppression essentially survives, as Figs.4(c,d) and 5(b) show. Moreover, its magnitude remains close to that in the 1D case. Thirdly, in the 2D ion with the interacting electrons (Fig. 4(c)), the diagonal flux, although much suppressed compared to the case of non-interactive electrons (Fig.4(d)), survives highly exceeding that of Fig.1(c). This flux may be caused by the collective tunneling of the electrons.
Figure 3: Time-dependent populations of singly (solid black lines) and doubly (dashed black lines) ionized atoms in the LPR (a) and SPR (b). The lower inserts show the electric field envelope.
Figure 4: Same as in Fig.1, but in the short pulse regime. The field parameters and the time instants are: $t=2.88$ fs (a), $t=2.16$ fs (b), $E_{0}=0.035$, $T=2$ fs.
Figure 5: Same as in Fig.2, but in the short pulse regime. Red lines show the total probability of double ionization (8) for the case of the interacting electrons, and blue lines – without the interaction. The field parameters are: $E_{0}=0.035$, $T=2$ fs.
In conclusion, we performed numerical TDSE solutions for a model two-electron negative ion with ionization potentials equal to those of ${\rm Br}^{-}$, subject to a strong time-dependent electric field in 1D and 2D space. Conditions of the tunnel field ionization were fulfilled for both the outer and inner electrons. The aim was to find to what extent and under which conditions the e-e interaction influences the rate of double ionization destroying the customary picture of sequential single-electron tunneling. Our results show that this happens indeed in short pulses, when the tunnel ionization rate for the inner electron becomes sizable before the outer electron is removed with guarantee. In this case, the probability of double ionization is suppressed by more than one order of magnitude, compared to the case of independent sequential tunneling. From the other side, our results show that in long pulses the sequential tunneling concept applies surprisingly well. Finally, we demonstrated weak but clear signatures of collective tunneling, which is almost entirely blocked by e-e repulsion in 1D, but survives in 2D and therefore should be expected even stronger in real 3D atoms and ions.
Our analysis fundamentally revisits the concept of collective tunneling in atoms introduced in [22, 23] 25 years ago. The results reported here unambiguously demonstrate that the semiclassical model [22] describing simultaneous tunneling of two electrons relying on the full discard of e-e repulsion greatly overestimates the probability: for the parameters chosen in our calculations this model would predict the near-diagonal flux comparable or even dominating over the sequential one. Also, our results explain why the collective tunneling contribution has not been detected in 1D TDSE numerical solutions of [23, 25]. Our results show also that the single active electron approximation routinely applied in the theory of strong field processes, particularly in the tunneling regime, is strongly limited for intense ultrashort laser pulses.
Our predictions may have an important impact at least in two research areas of strong-field atomic physics. Firstly, these results suggest the way to a search for collective tunneling in atomic systems. To this end, one needs to examine correlated momentum distributions of electron pairs for double ionization in a circularly polarized laser pulse, which acts efficiently as a unipolar pulse used in the present calculations. The qualitative difference between the density distributions of Figs.1(c) and 4(c) should lead to the corresponding difference in the momentum distributions, which can be measured using the COLTRIMS technique [36]. Secondly, the predicted breakdown of the sequential tunneling picture should be observable and even playing a quantitatively important role in the process of multiple ionization of atoms in short laser pulses of extreme intensities, where, due to the high peak field and the steepness of the time envelope the saturation time for $d$ and $p$ sub-shells can lie in the sub-fs domain. Taking into account that relative values of the ionization potentials within such sub-shells vary from one level to another much less than in the considered negative bromine ion, one should expect even stronger deviations from the sequential ionization picture than those demonstrated here. In particular, the suppression of ionization induced by the e-e correlation should be of interest for the development of numerical particle-in-cell codes aimed at modeling of laser-plasma dynamics with multiple field ionization of atoms incorporated [37].
DIT and SVP acknowledge support of the Theoretical Physics and Mathematics Advancement Foundation “BASIS”.